Two point charges experience an attractive force of 8 N when they are separated by 1 m. What is the magnitude of the force they experience when their separation is 0.5 m?
Answer #1

electrostatic force is inversely proportional to the square of distance,hence,

F2/F1 = (R1/R2)2

F2=F1 *(R1/R2)2

F2= 8*(1/0.5)2

F2= 32 N
Answer #2

SOLUTION :


F is inversely proportional to  squared distance, d^2

=> F = k / d^2 

=> k = F d^2

When F = 8 N and d = 1 m :

=> k = 8 * (1)^2 = 8 N-m^2


So, when distance is 0.5 m :


F2 = k  / (0.5)^2 = 8/ (0.5)^2 = 32 N (ANSWER).

answered by: Tulsiram Garg